# Resilience for Regular Path Queries: Towards a Complexity Classification

### v2

@@ Abstract.
We show that computing the resilience in this sense is tractable (even in combined complexity) for all RPQs defined from so-called local languages. By contrast, we show hardness in data complexity for RPQs defined from the following language classes (after reducing the languages to eliminate redundant words): all finite languages featuring a word containing a repeated letter, and all languages featuring a specific kind of counterexample to being local (which we call four-legged languages). The latter include in particular all languages that are not star-free. Our results also imply hardness for all non-local languages with a so-called neutral letter. We last highlight some remaining obstacles towards a full dichotomy.
@@ 1. Introduction
Figure 1. Overview of our results on the complexity of resilience (all results apply both to bag and set semantics), with some examples of languages. All languages are reduced. For legibility, we omit Section 5.6.
@@
The resilience problem has been introduced for Boolean conjunctive queries and mostly studied for such queries (DBLP:journals/pvldb/FreireGIM15; DBLP:conf/pods/FreireGIM20; bodirsky2024complexity; makhija2024unified). The key challenge in this line of work is to understand the data complexity of the problem, i.e., its complexity for a fixed query $Q$, as a function of the input database $D$. More precisely, the resilience problem can be tractably solved for some queries (e.g., trivial queries), and is known to be NP-hard for some other queries (see, e.g., (makhija2024unified)). Membership in NP always holds whenever query evaluation for $Q$ can be done in polynomial data complexity.
@@
In this paper, we study the resilience problem from a different perspective: we focus on graph databases (i.e., databases on an arity-two signature), and on queries which are defined by regular languages, namely, Boolean Regular Path Queries (Boolean RPQs), with no constants and with existentially quantified endpoints. In other words, the query asks whether the database contains a walk whose edge labels forms a word of the regular language, and resilience asks how many database facts must be removed so that no such walk exists. To our knowledge, the only other work which studies this problem is the recent paper of Bodirsky et al. (bodirsky2024complexity), in which the authors use CSP techniques. Their results apply to much more general languages than CQs, and in particular imply a dichotomy on the resilience problem for Boolean RPQs (and also for the broader class of so-called two-way RPQs aka 2RPQs). However, their lower bounds do not apply to databases in set semantics. In addition, their dichotomy criterion, while decidable, does not give a intuitive language-theoretic characterization of the tractable cases. We discuss this in more details at the end of Section 2.
@@
Our study of resilience for RPQs is motivated by database applications, but it is also motivated by connections to classical network flow problems. Consider for instance the classical MinCut problem of finding the minimum-cost cut in a flow network. The MinCut problem, allowing multiple sources and sinks without loss of generality, is easily seen to be equivalent to the resilience problem in bag semantics for the RPQ $ax^{*}b$ on the alphabet $\{a,b,x\}$, with $a$-facts of the database corresponding to sources, $b$-facts to sinks, and $x$-facts to edges of the network. Hence, the resilience problem for RPQs corresponds to a generalization of MinCut, where the edges of the flow network are labeled and we must cut edges to ensure that no walk in a specific language remains. In the other direction, for many RPQs, we can devise tractable algorithms for resilience by reducing to a MinCut problem on a suitably-constructed flow network. However, resilience is NP-hard for some RPQs, e.g., the finite RPQ $aa$ (DBLP:conf/pods/FreireGIM20), so these methods do not always apply. Our results aim at determining what are the RPQs for which resilience can be tractably computed (in particular via such methods), and what are the RPQs for which it is intractable.
@@
After giving preliminary definitions and stating the problem in Section 2, we start in Section 3 by giving our main tractability result (Theorem 3.3): resilience is tractable for RPQs defined from so-called local languages (mpri; sakarovitch2009elements). Those languages are intuitively those that are defined by specifying the possible initial labels of a path, the possible final labels of a path, and the consecutive pairs of labels that are allowed. We show that, for RPQs defined with a local language $L$, we can compute resilience by reducing to the MinCut problem via a natural product construction. The construction applies to queries both in set and bag semantics, and it is also tractable in combined complexity, running in time $\tilde{O}(|A|\times|\Sigma|\times|D|)$ where $A$ is an automaton representing $L$, where $\Sigma$ is the signature, and where $D$ is the input database.
@@
We then turn to intractability results in Sections 4, 5 and 6. In all these results, we assume that the language has been reduced to eliminate redundant words, e.g., resilience for $abbc|bb$ and $bb$ is the same because the corresponding (existentially-quantified) Boolean RPQs are equivalent.
@@
We define in Section 4 the notion of a hardness gadget for a language $L$, namely, a specific database to be used as a building block in a reduction. These gadgets are the analogue, in our setting, of the Independent Join Paths (IJPs) used in earlier work (makhija2024unified) to show hardness for CQs. When an RPQ $Q$ has a hardness gadget $\Gamma$, we can reduce the minimum vertex cover problem on an undirected graph $G$ to a resilience computation (in set semantics) for $Q$ on a database built using $\Gamma$, by replacing the edges of $G$ by a copy of $\Gamma$. We illustrate in Section 4 how our gadgets can be used to show hardness for two specific RPQs: $axb|cxd$ (Proposition 4.8), and $aa$ (Proposition 4.1).
@@
We then generalize these two hardness results in turn to apply to a broader class of languages. First, in Section 5, we introduce the notion of a four-legged language, as a language featuring a specific kind of counterexample to being local. More precisely, the local languages of Section 3 have the following known equivalent characterization (sakarovitch2009elements): for any letter $x$ and words $\alpha,\beta,\gamma,\delta$, if a language $L$ contains the words $\alpha x\beta$ and $\gamma x\delta$ but does not contain the “cross-product” word $\alpha x\delta$, then $L$ is not local. A four-legged language is a non-local language that features a counterexample of this form in which all the words $\alpha,\beta,\gamma,\delta$ are required to be non-empty. This covers, for instance, $axb|cxd$, but not $ab|bc$. We show (Section 5.2) that resilience is intractable for any four-legged language. This result implies hardness for many languages: it covers all non-star-free languages (Section 5.5), a well-known class of infinite regular languages; and, along with the tractability of local languages, it implies a dichotomy for languages having a neutral letter (Section 5.6).
@@
Second, in Section 6, we show that resilience is hard for any finite reduced language containing a word with a repeated letter (Section 6.1). This generalizes the case of $aa$, and shows that self-joins always imply hardness in the following sense: finite RPQs (those equivalent to UCQs), once reduced, have intractable resilience as soon as one of their constituent CQs features a self-join. This contrasts with the setting of CQs, where characterizing the tractable queries with self-joins is still open.
@@
Note that our two hardness results (Section 5.2 and Section 6.1) are incomparable: languages with a repeated letter are non-local, but they are not all four-legged (e.g., $aa$); conversely, some four-legged languages have no words with a repeated letter (e.g., $axb|cxd$). Assuming finiteness in Section 6.1 is also crucial, as $ax^{*}b$ features words with repeated letters and still enjoys tractable resilience. The proof of Section 6.1 is our main technical achievement: the proof relies on Section 5.2 and uses different gadgets depending on which additional words the language contains.
@@
Section 7 explores how our results do not classify all RPQs. We focus on finite RPQs: the remaining open cases are then the RPQs which have no words with repeated letters, are not local, but are not four-legged. We introduce the class of chain queries (e.g., $axb|byc$) to cover some of these RPQs, and show tractability (and reducibility to MinCut) assuming a bipartiteness condition. We also study the RPQ $abc|be$ and variants thereof, and show that it enjoys tractable resilience via techniques from submodular optimization (Section 7.4), which seems new in this context. This query (i.e., $R(x,y),S(y,z),T(z,u)\vee S(x,y),U(y,z)$ as a UCQ), and variants thereof, are currently the only known tractable RPQs with no known reduction to MinCut. We conclude by showing that some other unclassified RPQs are hard, and by discussing some remaining open cases.
We finally conclude in Section 8. For lack of space, most proofs are deferred to the appendix.
@@ 2. Preliminaries and Problem Statement
A language $L$ is a (possibly infinite) subset of $\Sigma^{*}$. For $L,L^{\prime}$ two languages, we write their concatenation as $LL^{\prime}\coloneq\{\alpha\beta\mid\alpha\in L,\beta\in L^{\prime}\}$. We mainly focus in this work on regular languages, also known as rational languages, which can be represented by finite state automata. We use the following automaton formalisms, which recognize precisely the regular languages (sakarovitch2009elements): deterministic finite automata (DFA), nondeterministic finite automata (NFA), and NFAs with $\epsilon$-transitions ($\epsilon$NFA, also called automata with spontaneous transitions).
An $\epsilon$NFA $A=(S,I,F,\Delta)$ is a finite set $S$ of states, a set $I\subseteq S$ of initial states, a set $F\subseteq S$ of final states, and a transition relation $\Delta\subseteq S\times(\Sigma\cup\{\epsilon\})\times S$. Let $\alpha=a_{1}\cdots a_{m}$ be a word. A configuration for $A$ on $\alpha$ is an ordered pair $(s,i)$ for $s\in S$ and $0\leq i\leq m$. A run of $A$ on $\alpha$ is a sequence of configurations $(s_{0},i_{0}),\ldots,(s_{\ell},i_{\ell})$ with $\ell\geq m$ such that $s_{0}\in I$ is initial, $i_{0}=0$, $i_{\ell}=m$, and for each $0\leq j<\ell$ we have either $i_{j+1}=i_{j}$ and $(s_{j},\epsilon,s_{j+1})\in\Delta$ (an $\epsilon$-transition) or $i_{j+1}=i_{j}+1$ and $(s_{j},a_{i_{j}+1},s_{j+1})\in\Delta$ (a letter transition). The run is accepting if $s_{\ell}\in F$. The language of $A$, denoted $\mathcal{L}(A)$, is the set of words having an accepting run: we say that $A$ recognizes $\mathcal{L}(A)$. An example $\epsilon$NFA $A_{3}$ is depicted in Figure 2(c), and an example accepting run of $A_{3}$ on the word $ad$ is $(s_{1},0),(s_{2},1),(s_{4},1),(s_{5},2)$. The size $|A|$ of the $\epsilon$NFA $A=(S,I,F,\Delta)$ is $|S|+|\Delta|$, its total number of states and transitions.
An NFA is an $\epsilon$NFA which does not feature $\epsilon$-transitions, and a DFA is an NFA having exactly one initial state and having at most one transition $(s,a,s^{\prime})$ in $\Delta$ for each $s\in S$ and $a\in\Sigma$.
We also use regular expressions to denote regular languages: we omit their formal definition, see (sakarovitch2009elements). We write |||| for the union operator and * for the Kleene star operator.
@@
[…]
@@
Let $D$ be a database and $Q$ a Boolean query. The resilience of $Q$ on $D$, denoted $\mathtt{RES}_{\mathrm{set}}(Q,D)$, is defined [^1] By convention, if $Q$ is satisfied by all subsets of $D$, then the minimum is defined to be +$\infty$+ ∞. as $\mathtt{RES}_{\mathrm{set}}(Q,D)\coloneq\min_{\begin{subarray}{c}D^{\prime}\subseteq D\\ \text{s.t. }D\setminus D^{\prime}\not\models Q\end{subarray}}|D^{\prime}|$.
In other words, the resilience is the minimum number of facts to remove from $D$ so that $Q$ is not satisfied. A subset of facts $D^{\prime}\subseteq D$ such that $D\setminus D^{\prime}\not\models Q$ is called a cut (or a contingency set) of $D$ for $Q$. Notice that if $D\not\models Q$, then $\emptyset$ is a cut and we have [^2] We deviate from the original definition in (DBLP:journals/pvldb/FreireGIM15) which left resilience undefined for the case of $D\not\models Q$. $\mathtt{RES}_{\mathrm{set}}(Q,D)=0$.
@@
From a language $L$ over $\Sigma$ we now define a query over graph databases that tests the existence of walks labeled by words of $L$. Given a database $D$, a walk in $D$ is a (possibly empty) sequence of consecutive edges $v_{1}\xrightarrow{a_{1}}v_{2}$, $v_{2}\xrightarrow{a_{2}}v_{3}$, $\ldots$, $v_{k}\xrightarrow{a_{k}}v_{k+1}$ in $D$. Note that the $a_{i}$ and $v_{i}$ are not necessarily pairwise distinct. We say that the walk is labeled by the word $\alpha=a_{1}\cdots a_{k}$, and call it an $\alpha$-walk. We also say that the walk contains a word $\beta\in\Sigma^{*}$ if $\beta$ is an infix of $\alpha$. An $L$-walk in $D$ is then an $\alpha$-walk for some $\alpha\in L$. We define the Boolean query $Q_{L}$ as: $Q_{L}(D)=1$ iff there is some $L$-walk in $D$. In particular, if $\epsilon\in L$ then $Q_{L}(D)=1$ on every database $D$ (including $D=\emptyset$). A regular path query (RPQ) is then a Boolean query of the form $Q_{L}$ for $L$ a regular language.
Observe that $Q_{L}$ does not change if we remove from $L$ all words $\beta$ such that some strict infix $\alpha$ of $\beta$ is also in $L$: indeed if $D$ contains a $\beta$-walk then it also contains an $\alpha$-walk. Formally, the reduced language of $L$ is $\mathtt{reduce}(L)\coloneq\{\alpha\in L\mid\text{no strict infix of }\alpha\text{ is in }L\}$. We then have that $Q_{L}$ and $Q_{\mathtt{reduce}(L)}$ are the exact same queries. The reduced language of $L$ is also called the infix-free sublanguage of $L$ in (barcelo2019boundedness): if $L$ is regular then so is $\mathtt{reduce}(L)$. We say that $L$ is reduced if $L=\mathtt{reduce}(L)$.
@@
We also consider the resilience problems on bag graph databases, by opposition with set graph databases which are the databases defined so far. A bag (graph) database consists of a database $D$ together with a multiplicity function $\mathtt{mult}:D\to\mathbb{N}_{>0}$; it can equivalently be seen as a multiset of facts. In the input, the multiplicities are expressed as integers written in binary. A Boolean query $Q$ is satisfied by a bag database $(D,\mathtt{mult})$ iff $D$ satisfies $Q$, i.e., the multiplicities of facts are not seen by queries. A contingency set for $Q$ and $D$ is then a subset $D^{\prime}\subseteq D$ such that $D\setminus D^{\prime}\not\models Q$, and the resilience of $Q$ on $D$ is: $\mathtt{RES}_{\mathrm{bag}}(Q,D)\coloneq\min_{\begin{subarray}{c}D^{\prime}\subseteq D\\ \text{s.t. }D\setminus D^{\prime}\not\models Q\end{subarray}}\sum_{e\in D^{\prime}}\mathtt{mult}(e)$, i.e., fact multiplicities represent costs.
@@
Most of our resilience algorithms work by reducing to the problem of finding a minimum cut of a flow network (Cormen:2009dz). Recall that a flow network consists of a directed graph $\mathcal{N}=(V,t_{\text{source}},t_{\text{target}},E,c)$ with $V$ being the vertices, $t_{\text{source}},t_{\text{target}}\in V$ being special vertices called the source and the target respectively, $E\subseteq V\times V$, and $c:E\to\mathbb{R}_{+}\cup\{+\infty\}$ giving the capacity of each edge. For $E^{\prime}\subseteq E$, we define the cost of $E^{\prime}$ as $c(E^{\prime})\coloneq\sum_{e\in E^{\prime}}c(e)$, and denote $\mathcal{N}\setminus E^{\prime}$ as the network in which all edges from $E^{\prime}$ are removed. A cut is a set of edges $E^{\prime}\subseteq E$ such that there is no directed path from $t_{\text{source}}$ to $t_{\text{target}}$ in $\mathcal{N}\setminus E^{\prime}$. The min-cut problem, written MinCut, takes as input a flow network and returns the minimum cost of a cut. From the max-flow min-cut theorem (ford1956maximal) and Menger’s theorem (menger:1927), we know that the min-cut problem can be solved in PTIME. More precisely, practical algorithms can run in $O(|V|^{2}\sqrt{|E|})$ (ahuja1997computational), and recent theoretical algorithms run in $\tilde{O}(|\mathcal{N}|)$ (henzinger2024deterministic), where $|\mathcal{N}|\coloneq|V|+|E|$ and where $\tilde{O}$ hides polylogarithmic factors.
@@
In this paper, we study the resilience problem for queries $Q_{L}$ defined by languages $L$; most of our results are for regular languages. The resilience problem for $L$, written $\mathtt{RES}_{\mathrm{set}}(L)$, is the following: given as input a database $D$ and $k\in\mathbb{N}$, decide if $\mathtt{RES}_{\mathrm{set}}(Q_{L},D)\leq k$. The problem $\mathtt{RES}_{\mathrm{bag}}(L)$ in bag semantics is defined analogously. Note that $\mathtt{RES}_{\mathrm{set}}(L)$ easily reduces to $\mathtt{RES}_{\mathrm{bag}}(L)$ by giving multiplicity of 1 to each fact; however, in the setting of CQs, some conjunctive queries $Q_{0}$ are known for which $\mathtt{RES}_{\mathrm{set}}(Q_{0})$ is in PTIME, but $\mathtt{RES}_{\mathrm{bag}}(Q_{0})$ is NP-hard (makhija2024unified, Table 1) (even when the multiplicities of the input bag database are written in unary).

### v3

@@ Abstract.
We show that computing the resilience in this sense is tractable (even in combined complexity) for all RPQs defined from so-called local languages. By contrast, we show hardness in data complexity for RPQs defined from the following language classes (after reducing the languages to eliminate redundant words): all finite languages featuring a word containing a repeated letter, and all languages featuring a specific kind of counterexample to being local (which we call four-legged languages). The latter include in particular all languages that are not star-free. Our results also imply hardness for all non-local languages with a so-called neutral letter. We last show tractability for some classes of non-local languages, namely the so-called bipartite chain languages and one-dangling languages, and highlight some remaining obstacles towards a full dichotomy.
@@ 1. Introduction
Figure 1. Overview of our results on the complexity of resilience (all results apply both to bag and set semantics), with some examples of languages. All languages are infix-free. For legibility, we omit Proposition 5.7.
@@
The resilience problem has been introduced for Boolean conjunctive queries and mostly studied for such queries (DBLP:journals/pvldb/FreireGIM15, ; DBLP:conf/pods/FreireGIM20, ; bodirsky2024complexity, ; makhija2024unified, ). The key challenge in this line of work is to understand the data complexity of the problem, i.e., its complexity for a fixed query $Q$, as a function of the input database $D$. More precisely, the resilience problem can be tractably solved for some queries (e.g., trivial queries), and is known to be NP-hard for some other queries (see, e.g., (makhija2024unified, )). Membership in NP always holds whenever query evaluation for $Q$ can be done in polynomial data complexity.
@@
In this paper, we study the resilience problem from a different perspective: we focus on graph databases (i.e., databases on an arity-two signature), and on queries which are defined by regular languages, namely, Boolean Regular Path Queries (Boolean RPQs), with no constants and with existentially quantified endpoints. In other words, the query asks whether the database contains a walk whose edge labels forms a word of the regular language, and resilience asks how many database facts must be removed so that no such walk exists. To our knowledge, the only other work which studies this problem is the recent paper of Bodirsky et al. (bodirsky2024complexity, ), in which the authors use CSP techniques. Their results apply to much more general languages than CQs, and in particular imply a dichotomy on the resilience problem for Boolean RPQs (and also for the broader class of so-called two-way RPQs aka 2RPQs). However, their lower bounds do not apply to databases in set semantics. In addition, their dichotomy criterion, while decidable, does not give an intuitive language-theoretic characterization of the tractable cases. We discuss this in more detail at the end of Section 2.
@@
Our study of resilience for RPQs is motivated by database applications, but it is also motivated by connections to classical network flow problems. Consider for instance the classical MinCut problem of finding the minimum-cost cut in a flow network. The MinCut problem, allowing multiple sources and sinks without loss of generality, is easily seen to be equivalent to the resilience problem in bag semantics for the RPQ $ax^{*}b$ on the alphabet $\{a,b,x\}$, with $a$-facts of the database corresponding to sources, $b$-facts to sinks, and $x$-facts to edges of the network. Hence, the resilience problem for RPQs corresponds to a generalization of MinCut, where the edges of the flow network are labeled and we must cut edges to ensure that no walk in a specific language remains. In the other direction, for many RPQs, we can devise tractable algorithms for resilience by reducing to a MinCut problem on a suitably-constructed flow network. However, resilience is NP-hard for some RPQs, e.g., the finite RPQ $aa$ (DBLP:conf/pods/FreireGIM20, ), so these methods do not always apply. Our results aim at determining what are the RPQs for which resilience can be tractably computed (in particular via such methods), and what are the RPQs for which it is intractable.
@@
After giving preliminary definitions and stating the problem in Section 2, we start in Section 3 by giving our main tractability result (Theorem 3.13): resilience is tractable for RPQs defined from so-called local languages (mpri, ; sakarovitch2009elements, ). Those languages are intuitively those that are defined by specifying the possible initial labels of a path, the possible final labels of a path, and the consecutive pairs of labels that are allowed. We show that, for RPQs defined with a local language $L$, we can compute resilience by reducing to the MinCut problem via a natural product construction. The construction applies to queries both in set and bag semantics, and it is also tractable in combined complexity, running in time $\tilde{O}(|A|\times|\Sigma|\times|D|)$ where $A$ is an automaton representing $L$, where $\Sigma$ is the signature, and where $D$ is the input database.
@@
We then turn to intractability results in Sections 4, 5 and 6. In all these results, we assume that we focus on the infix-free sublanguage where redundant words have been eliminated, e.g., resilience for $abbc|bb$ and $bb$ is the same because the corresponding (existentially-quantified) Boolean RPQs are equivalent.
@@
We define in Section 4 the notion of a hardness gadget for a language $L$, namely, a specific database to be used as a building block in a reduction. These gadgets are the analogue, in our setting, of the Independent Join Paths (IJPs) used in earlier work (makhija2024unified, ) to show hardness for CQs. When an RPQ $Q$ has a hardness gadget $\Gamma$, we can reduce the minimum vertex cover problem on an undirected graph $G$ to a resilience computation (in set semantics) for $Q$ on a database built using $\Gamma$, by replacing the edges of $G$ by a copy of $\Gamma$. We illustrate in Section 4 how our gadgets can be used to show hardness for two specific RPQs: $axb|cxd$ (Proposition 4.13), and $aa$ (Proposition 4.1).
@@
We then generalize these two hardness results in turn to apply to a broader class of languages. First, in Section 5, we introduce the notion of a four-legged language, as a language featuring a specific kind of counterexample to being local. More precisely, the local languages of Section 3 have the following known equivalent characterization (sakarovitch2009elements, ): for any letter $x$ and words $\alpha,\beta,\gamma,\delta$, if a language $L$ contains the words $\alpha x\beta$ and $\gamma x\delta$ but does not contain the “cross-product” word $\alpha x\delta$, then $L$ is not local. A four-legged language is a non-local language that features a counterexample of this form in which all the words $\alpha,\beta,\gamma,\delta$ are required to be non-empty. This covers, for instance, $axb|cxd$, but not $ab|bc$. We show (Theorem 5.3) that resilience is intractable for any four-legged language. This result implies hardness for many languages: it covers all non-star-free languages (Lemma 5.6), a well-known class of infinite regular languages; and, along with the tractability of local languages, it implies a dichotomy for languages having a neutral letter (Proposition 5.7).
@@
Second, in Section 6, we show that resilience is hard for any finite infix-free language containing a word with a repeated letter (Theorem 6.1). This generalizes the case of $aa$, and shows that self-joins always imply hardness in the following sense: finite RPQs (those equivalent to UCQs), once made infix-free, have intractable resilience as soon as one of their constituent CQs features a self-join. This contrasts with the setting of CQs, where characterizing the tractable queries with self-joins is still open.
@@
Note that our two hardness results (Theorem 5.3 and Theorem 6.1) are incomparable: languages with a repeated letter are non-local, but they are not all four-legged (e.g., $aa$); conversely, some four-legged languages have no words with a repeated letter (e.g., $axb|cxd$). Assuming finiteness in Theorem 6.1 is also crucial, as $ax^{*}b$ features words with repeated letters and still enjoys tractable resilience. The proof of Theorem 6.1 is our main technical achievement: it relies on Theorem 5.3 and uses different gadgets depending on which additional words the language contains.
@@
Section 7 then explores how our results do not classify all RPQs. We mostly focus on finite RPQs: the remaining open cases are then the RPQs which have no words with repeated letters, are not local, but are not four-legged. We introduce the class of chain languages (e.g., $axb|byc$) to cover some of these RPQs, and show tractability (and reducibility to MinCut) assuming a bipartiteness condition. We then introduce the class of one-dangling languages (e.g., $abc|be$), for which we show tractability (via a more elaborate reducibility to MinCut). We finish the section by showing that some other unclassified RPQs are intractable. We conclude the paper in Section 8.
Conference version. This paper is an extended version of an earlier conference paper (amarilli2025resilience, ). Compared to the conference version, all our tractability results in the current paper use a unified reduction to flow. Some of the new proofs are considerably simpler, and we currently do not have any example of a query with tractable resilience for which tractability does not follow by reducing to an appropriate network flow problem. Our current results support our hypothesis that flow is all one needs for tractable cases, even those that are currently open. In more detail, we have improved Proposition 7.9 (compared to (amarilli2025resilience, , Prp. 7.7)) by showing a tractability result that covers a larger family of languages (dubbed one-dangling languages) and uses a reduction to a flow problem. This yields tractability in combined complexity and gives a better exponent than an earlier proof via submodular optimization from (amarilli2025resilience, ). An overview of our results in the present paper is in Figure 1. Compared with (amarilli2025resilience, , Figure 1), the languages $abcd|be$ and $ax^{*}b|xd$ have now been classified as tractable thanks to the new Proposition 7.9, and new unclassified languages $ab^{*}c|ba$ and $ab^{*}d|ac^{*}d|bc$ were added. The present paper also includes the complete proofs for all results, which had been omitted from the conference article.
@@ 2. Preliminaries and Problem Statement
[…]
We also use regular expressions to denote regular languages: we omit their formal definition, see (sakarovitch2009elements, ). We write $|$ for the union operator and * for the Kleene star operator.
@@
A graph database on the alphabet $\Sigma$, or simply database, is a set of $\Sigma$-labeled edges of the form $D\subseteq V\times\Sigma\times V$, where $V$ is the set of nodes of $D$: we also call $V$ the domain and write it as $\mathrm{Adom}(D)$. Notice that a graph database can equivalently be seen as a relational database on an arity-two signature that directly corresponds to the alphabet $\Sigma$. We also call facts the edges $(v,a,v^{\prime})$ of $D$, and write them as $v\xrightarrow{a}v^{\prime}$. We call $v$ the source (or tail) of $v\xrightarrow{a}v^{\prime}$ and $v^{\prime}$ its target (or head). A Boolean query is a function $Q$ that takes as input a database $D$ and outputs $Q(D)\in\{0,1\}$. We write $D\models Q$, and say that $Q$ is satisfied by $D$, when $Q(D)=1$, and write $D\not\models Q$ otherwise. In this paper we study the resilience of queries:
@@
Let $D$ be a database and $Q$ a Boolean query. The resilience of $Q$ on $D$, denoted $\mathtt{RES}_{\mathrm{set}}(Q,D)$, is defined [^1] By convention, if $Q$ is satisfied by all subsets of $D$, then the minimum is defined to be $\infty$. as $\mathtt{RES}_{\mathrm{set}}(Q,D)\coloneq\min_{\begin{subarray}{c}D^{\prime}\subseteq D\\ \text{s.t. }D\setminus D^{\prime}\not\models Q\end{subarray}}|D^{\prime}|$.
In other words, the resilience is the minimum number of facts to remove from $D$ so that $Q$ is not satisfied. A subset of facts $D^{\prime}\subseteq D$ such that $D\setminus D^{\prime}\not\models Q$ is called a cut (or a contingency set) of $D$ for $Q$. Notice that if $D\not\models Q$, then $\emptyset$ is a cut and we have [^2] We deviate from the original definition in (DBLP:journals/pvldb/FreireGIM15, ) which left resilience undefined for the case of $D\not\models Q$. $\mathtt{RES}_{\mathrm{set}}(Q,D)=0$.
@@
From a language $L$ over $\Sigma$ we now define a query over graph databases that tests the existence of walks labeled by words of $L$. Given a database $D$, a walk in $D$ is a (possibly empty) sequence of consecutive edges $v_{1}\xrightarrow{a_{1}}v_{2}$, $v_{2}\xrightarrow{a_{2}}v_{3}$, $\ldots$, $v_{k}\xrightarrow{a_{k}}v_{k+1}$ in $D$. Note that the $a_{i}$ and $v_{i}$ are not necessarily pairwise distinct. We say that the walk is labeled by the word $\alpha=a_{1}\cdots a_{k}$, and call it an $\alpha$-walk. We also say that the walk contains a word $\beta\in\Sigma^{*}$ if $\beta$ is an infix of $\alpha$. An $L$-walk in $D$ is then an $\alpha$-walk for some $\alpha\in L$. We define the Boolean query $Q_{L}$ as: $Q_{L}(D)=1$ iff there is some $L$-walk in $D$. In particular, if $\epsilon\in L$ then $Q_{L}(D)=1$ on every database $D$ (including $D=\emptyset$). A regular path query (RPQ) is then a Boolean query of the form $Q_{L}$ for $L$ a regular language. Note that, in all of the present paper, we evaluate RPQs with walk semantics (i.e., arbitrary paths with repetitions), as opposed to other works such as (martens2019dichotomies, ) which evaluate them with simple path or trail semantics: we leave open the question of studying resilience for RPQs in simple path or trail semantics.
Observe that $Q_{L}$ does not change if we remove from $L$ all words $\beta$ such that some strict infix $\alpha$ of $\beta$ is also in $L$: indeed if $D$ contains a $\beta$-walk then it also contains an $\alpha$-walk. Formally, the infix-free language of $L$ is $\mathtt{IF}(L)\coloneq\{\alpha\in L\mid\text{no strict infix of }\alpha\text{ is in }L\}$. We then have that $Q_{L}$ and $Q_{\mathtt{IF}(L)}$ are the exact same queries. The infix-free language of $L$ can also be called the factor-free sublanguage of $L$ in (barcelo2019boundedness, ), and is called the reduced sublanguage in our earlier conference paper (amarilli2025resilience, ). It is easy to show that if $L$ is regular then so is $\mathtt{IF}(L)$. We say that $L$ is infix-free if $L=\mathtt{IF}(L)$; this is also called being an infix code (ito1991outfix, ).
@@
We also consider the resilience problems on bag graph databases, as opposed to set graph databases which are the databases defined so far. A bag (graph) database consists of a database $D$ together with a multiplicity function $\mathtt{mult}:D\to\mathbb{N}_{>0}$; it can equivalently be seen as a multiset of facts. In the input, the multiplicities are expressed as integers written in binary. A Boolean query $Q$ is satisfied by a bag database $(D,\mathtt{mult})$ iff $D$ satisfies $Q$, i.e., the multiplicities of facts are not seen by queries. A contingency set of $D$ for $Q$ is then again a subset $D^{\prime}\subseteq D$ such that $D\setminus D^{\prime}\not\models Q$, and the resilience of $Q$ on $D$ is: $\mathtt{RES}_{\mathrm{bag}}(Q,D)\coloneq\min_{\begin{subarray}{c}D^{\prime}\subseteq D\\ \text{s.t. }D\setminus D^{\prime}\not\models Q\end{subarray}}\sum_{e\in D^{\prime}}\mathtt{mult}(e)$, i.e., fact multiplicities represent costs.
@@
Our resilience algorithms work by reducing to the problem of finding a minimum cut of a flow network (Cormen:2009dz, ). Recall that a flow network consists of a directed graph $\mathcal{N}=(V,t_{\text{source}},t_{\text{target}},E,c)$ with $V$ being the vertices, $t_{\text{source}},t_{\text{target}}\in V$ being special vertices called the source and the target respectively, $E\subseteq V\times V$, and $c:E\to\mathbb{R}_{+}\cup\{+\infty\}$ giving the capacity of each edge. For $E^{\prime}\subseteq E$, we define the cost of $E^{\prime}$ as $c(E^{\prime})\coloneq\sum_{e\in E^{\prime}}c(e)$, and denote $\mathcal{N}\setminus E^{\prime}$ as the network in which all edges from $E^{\prime}$ are removed. A cut is a set of edges $E^{\prime}\subseteq E$ such that there is no directed path from $t_{\text{source}}$ to $t_{\text{target}}$ in $\mathcal{N}\setminus E^{\prime}$. The min-cut problem, written MinCut, takes as input a flow network and returns the minimum cost of a cut. From the max-flow min-cut theorem (ford1956maximal, ) and Menger’s theorem (menger:1927, ), we know that the min-cut problem can be solved in PTIME. More precisely, practical algorithms can run in $O(|V|^{2}\sqrt{|E|})$ (ahuja1997computational, ), and recent theoretical algorithms run in $\tilde{O}(|\mathcal{N}|)$ (henzinger2024deterministic, ), where $|\mathcal{N}|\coloneq|V|+|E|$ and where $\tilde{O}$ hides polylogarithmic factors.
@@
In this paper, we study the resilience problem for queries $Q_{L}$ defined by languages $L$; most of our results are for regular languages. The resilience problem for $L$, written $\mathtt{RES}_{\mathrm{set}}(L)$, is the following: given as input a database $D$ and $k\in\mathbb{N}$, decide if $\mathtt{RES}_{\mathrm{set}}(Q_{L},D)\leq k$. The problem $\mathtt{RES}_{\mathrm{bag}}(L)$ in bag semantics is defined analogously. Note that $\mathtt{RES}_{\mathrm{set}}(L)$ easily reduces to $\mathtt{RES}_{\mathrm{bag}}(L)$ by giving multiplicity of 1 to each fact; however, in the setting of CQs, some conjunctive queries $Q_{0}$ are known for which $\mathtt{RES}_{\mathrm{set}}(Q_{0})$ is in PTIME, but $\mathtt{RES}_{\mathrm{bag}}(Q_{0})$ is NP-complete (makhija2024unified, , Table 1) (even when the multiplicities of the input bag database are written in unary). We also note that the definition given here corresponds to a decision problem of determining whether the resilience is no greater than the value $k$. However, thanks to self-reducibility, the function problem of computing a witness contingency set reduces in polynomial time to this decision problem: see (bodirsky2025complexity, , Remark 8.4) for a self-contained argument.